Suppose a local bank increases the fees they charge for their bank accounts by 10 percent. In response, the demand for their bank a to 10,000. What is price elasticity of demand for this banks accounts? ccounts decreases from 25,000 Using the midpoint formula, the price elasticity of demand is (Enter your response rounded to two decimal places.)
Answer #1

Mid point elasticity has change in quantity = (10000 - 25000)/((25000 + 10000)/2) = -85.71%. The % change in price is 10.

Hence price elasticity is ed = % change in Qd/ % change in price = -85.71/10 = -8.57. (You should enter a positive value in case a negative sign is not accepted).